campare eletrostatics force &gravtational force b/w two protons

Grade:12

1 Answers

Arun
6 years ago
Lets compare it, suppose E is electrostatic force and F is gravitational force
E/F=kq^2/Gm^2
where m and q are mass and charge of electron respectively.
E/F=(9*10^9)*(1.6*10^-19)^2/(6.67*10^(-11))*(9.31*10^(-31))^2=3.985*10^42
ratio is almost 4*10^42.